OpenStudy (zolock):

can u simplify 5/3x^2 y - 4/6xy^3. yes or no if no explain why.

OpenStudy (mathstudent55):

This? \(\dfrac{5}{3x^2y} - \dfrac{4}{6xy^3} \)

OpenStudy (zolock):

yes

OpenStudy (mathstudent55):

First, the second fraction can be reduced since both the numerator and denominator are divisible by 2.

OpenStudy (mathstudent55):

Second, the fractions can be subtracted by first finding a common denominator.

OpenStudy (blank ):

I simplified it to: \[\frac{ (5y^2-2x) }{ (3x^2y^3) }\]

OpenStudy (zolock):

thanks guys

OpenStudy (blank ):

ǝɯoɔlǝʍ ǝɹ,noʎ :)

OpenStudy (mathstudent55):

\(\dfrac{5}{3x^2y} - \dfrac{4}{6xy^3}\) \(= \dfrac{5}{3x^2y} - \dfrac{2}{3xy^3} \) \(= \dfrac{5y^2}{3x^2y^3} - \dfrac{2x}{3x^2y^3} \) \(= \dfrac{5y^2 - 2x}{3x^2y^3} \)

OpenStudy (blank ):

Thanks @mathstudent55

OpenStudy (zolock):

thanks mathstudent55

OpenStudy (mathstudent55):

yw